Suncoast Corporation’s bonds have a 18-year maturity, a 5.5% semiannual coupon, and a par value of $1,000. The going interest rate (rd) is 8.25%, based on semiannual compounding.What is the bond’s price?
Answer #1

For this question,

M = $1000, n = 18 * 2 = 36 semi-annual periods, C = 5.5% * $1000/2 = $27.50 (semi-annually), i = 8.25%/2 = 4.125% (semi-annual)

P = $511.10 + $233.36

P = $744.45